$SPAD/src/input d01anf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
e:EXPR FLOAT:=log(X)
f:ASP1(G):=retract e
a:SF:=1.0e-6
b:SF:=1.0
epsabs:SF:=0.0
epsrel:SF:=1.0e-4
omega:SF:=10*%pi
liw:=max(400,2)
lw:=max(2*liw,4)
result:=d01anf(a,b,omega,2,epsabs,epsrel,lw,liw,-1,f)

———-

References^{[1] nothing}

4